Use the set of vectors provided to compute

for the resultant of the combined vectors

that follow.

A (short arrow above)= 28 units, 40° N of E

B (arrow above)= 11 units, 50° W of S

C (short arrow above)= 35 units, 40° S of W

D (arrow above)= 52 units, 50° E of N


1
Expert's answer
2021-10-26T17:18:04-0400
"R_x=28\cos{40}-11\sin{50}\\-35\cos{40}+52\sin{50}=26\\
R_y=28\sin{40}-11\cos{50}\\-35\sin{40}+52\cos{50}=22\\R=\sqrt{22^2+26^2}=34\\\theta=\arctan{\frac{22}{26}}=40\degree\text{ N of E}"